OpenStudy (anonymous):

Help I'm stuck on this question x/4 -5=3

OpenStudy (anonymous):

I got x/4=8

OpenStudy (eric_d):



OpenStudy (imstuck):

\[\frac{ x }{ 4 }-5=3\]Is that it?

OpenStudy (eric_d):

yea , carry on

OpenStudy (anonymous):

Yes

OpenStudy (eric_d):

x/4=8, incomplete

OpenStudy (anonymous):

I dnt kno what else to do it confusing

OpenStudy (eric_d):